solve triangelABC if A=24 degrees,B=62 degrees,and a= 21.4 (law of Sines
Found 2 solutions by solver91311, Alan3354:
Answer by solver91311(24713)



Now there is an interesting metaphysical phenomenon -- a triangel. One wonders how many triangels can dance on the head of a pin. Is it one third the number of regular angels?

But, taking a wild guess and presuming you really meant triangle ABC, let us proceed.

Law of Sines



We know angle A and angle B, so we can calculate angle C by remembering that the sum of the interior angles of a triangle is 180 degrees, hence:



And we know the measure of side a, so:



The rest is just punching calculator buttons.

John


Answer by Alan3354(69443)

solve triangelABC if A=24 degrees,B=62 degrees,and a= 21.4 (law of Sines)
--------------------
C = 94º
a/sin(A) = b/sin(B)
b = a*sin(B)/sin(A) =~ 46.455314
--------------------------------
Do the same for side c